OpenStudy (anonymous):

Solve: -2(4x+1)=5(-3x-2) Help

OpenStudy (anonymous):

Start by Distributing, Then Isolate X

OpenStudy (anonymous):

the answers are A. x=-3/7 B. x=12/23 C. x=-8/7 D. x=0

OpenStudy (anonymous):

\[-8x-2 = -15x -10\]\[7x = -8\]\[x = -8/7\]So, C

OpenStudy (anonymous):

Thank you so much I have a couple more if you can help ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve: 3(2x-4)=x+3-5x Answers: A. x=7/10 B. x=1 C. x=5 D. x=3/2

OpenStudy (anonymous):

\[6x-12 = -4x+3\]\[10x=15\]\[x=3/2\]So, D

OpenStudy (anonymous):

Thank you .

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I haveone more if u don't mind lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve: 3x-5=2(x+7) Answers: A. x=12 B. x=9 C. x=19 D. x=9/5

OpenStudy (anonymous):

is it hard ?